OpenStudy (anonymous):

The perimeter of a rectangle is 168 ft. The ratio of the length to the width is 11:1. Find the width of the rectangle. A. 7 ft B. 12 ft C. 14 ft D. 77 ft

OpenStudy (anonymous):

Let width be x Length = 11x Perimeter = 168 2(11x + x) = 168 2(12x) = 168 x = 168/24 = 7 Therefore, Length = 11x = 77 ft